# Capture of field stars by dark substructures

### v1

@@ Abstract
We use analytical and $N$-body methods to study the capture of field stars by gravitating substructures moving across a galactic environment. We find that the majority of captured stars move on temporarily-bound orbits that escape from the substructure potential after a few orbital revolutions. In numerical experiments where a substructure model is immersed in a galaxy on a circular orbit, we also find particles that remain bound to the substructure potential for indefinitely-long times. This population is absent from substructure models initially placed outside the galaxy on an eccentric orbit. We show that gravitational capture is most efficient in dwarf spheroidal galaxies (dSphs) on account of their low velocity dispersions and high stellar phase-space densities. In these galaxies ‘dark’ sub-subhaloes which do not experience in-situ star formation may capture field stars and become visible as stellar overdensities with unusual properties: (i) they would have a large size for their luminosity, (ii) contain stellar populations indistinguishable from the host galaxy, and (iii) exhibit dark matter (DM)-dominated mass-to-light ratios. We discuss the nature of several ‘anomalous’ stellar systems reported as star clusters in the Fornax and Eridanus II dSphs which exhibit some of these characteristics. A large population of DM sub-subhaloes with a mass function ${\rm d}N/{\rm d}M_{\bullet}\sim M_{\bullet}^{-\alpha}$ generates stellar systems with a luminosity function, ${\rm d}N/{\rm d}M_{\star}\sim M_{\star}^{-\beta}$, where $\beta=(2\alpha+1)/3=1.6$ for $\alpha=1.9$. Detecting and characterizing these objects in dSphs would provide unprecedented constraints on the particle mass and cross section of a large range of DM particle candidates.
@@ 2 Statistical theory
Section 2.1 follows the steps outlined in Paper I to compute the average number of field particles that have negative binding energies $E=v^{2}/2+\Phi_{\bullet}(r)<0$ within a spherical volume $V=4\pi r^{3}/3$ ($N_{b})$, where $r$ and $v$ are measured relative to the substructure, as well as their number density ($n_{\star}$) and velocity dispersion ($\sigma_{\star}$) profiles in dynamical equilibrium within a generic potential $\Phi_{\bullet}$. Here, we will focus on capture of field stars, although this analysis can be extended to other tracers moving in the host galaxy potential, including gas particles [^4] To model the dynamics of gaseous particles, a pressure term must be included in the equations of motion presented in §3.1.
@@ 2.1.1 Weak perturbations
The number of field stars temporarily bound to a moving substructure sourcing a potential $\Phi_{\bullet}$ can be estimated statistically from the (local) distribution function under the following simplifying assumptions: (i) field stars move on uncorrelated (random) trajectories within a small volume element $V$, (ii) their number density is roughly constant within the volume $V$, such that $n({\textbf{{R}}}_{\bullet}+{\textbf{{r}}})\approx n({\textbf{{R}}}_{\bullet})=n=N/V$, where ${\textbf{{R}}}_{\bullet}$ is the 3D position vector of the substructure within the host galaxy. This is known as the local approximation, and holds insofar as the (local) density profile rolls slowly, i.e. $r\ll d=|\nabla n/n|^{-1}$. (iii) The relative velocity distribution of particles within the volume element $V=4\pi r^{3}/3$ follows a Maxwellian distribution displaced by the reflex velocity of the point-mass $p(\mbox{\boldmath$v$})=(2\pi\sigma^{2})^{-3/2}\exp[-(\mbox{\boldmath$v$}+\mbox{\boldmath$V$}_{\bullet})^{2}/(2\sigma^{2})$ where $\sigma=\sigma(\mbox{\boldmath$R$}_{\bullet})$ is the local, one-dimensional velocity dispersion of field stars. Under the Maxwellian approximation, the mean-squared (relative) velocity between the background particles and the substructure is $\langle v^{2}\rangle=3\sigma^{2}+V_{\bullet}^{2}$, where ${\textbf{{V}}}_{\bullet}$ is the velocity vector of the substructure with respect to the host galaxy centre.
The approximation that field stars move on uncorrelated trajectories is accurate insofar as the number of stars bound to $\Phi_{\bullet}$ represents a small fraction of the total number of stars within the volume element, $N=n\,V$. Paper I shows that this condition requires an escape velocity $v_{e}\lesssim\sigma$, where $v_{e}(\mbox{\boldmath$r$})=\sqrt{2|\Phi_{\bullet}(\mbox{\boldmath$r$})|}$ is calculated at a distance $r=|\mbox{\boldmath$R$}-\mbox{\boldmath$R$}_{\bullet}|$ from the centre of the potential $\Phi_{\bullet}$.
@@
It is worth emphasizing that captured stars move on chaotic orbits that do not conserve energy or angular momentum, as shown in §3.2. However, from a statistical point of view, we can count how many stars are bound to the substructure potential (i.e. $E<0$) at any given time and compare that number against the value estimated in Equation (2). Hénon & Petit (1986) found that chaotic orbits only remain trapped within the potential $\Phi_{\bullet}$ over a finite amount of time before being lost to galactic tides. As shown in Paper I, this leads to a population of bound field stars that reaches a steady state as the capture rate (defined as the net number of field particles with $E$ flipping from positive to negative values) equals the loss rate (i.e the net number of field particles with $E$ flipping from negative to positive values). Paper I finds that steady state is typically reached over Smoluchowski’s (1916) ‘fluctuation mean life’
@@
The density enhancement due to the presence of trapped interstellar stars in steady state can be directly derived from (2) as
$\delta_{\star}(r)$ $\equiv\frac{1}{4\pi\,n\,r^{2}}\frac{{\rm d}N_{\rm ss}}{{\rm d}r}$ (4)
@@
$=\alpha\frac{4}{3\sqrt{\pi}}\frac{(GM_{\bullet})^{3/2}}{\sigma^{3}}e^{-V_{\bullet}^{2}/(2\sigma^{2})}\frac{1}{(r+c_{\bullet})^{3/2}},$
which converges to a centrally-divergent ‘density spike’ $\delta_{\star}\sim r^{-3/2}$ (Gondolo & Silk 1999) in the point-mass limit $c_{\bullet}\to 0$. For extended halos ($c_{\bullet}>0$) the density profile is flat for $r\lesssim c_{\bullet}$. Note that the density enhancement (4) scales as $\delta_{\star}\sim|\Phi_{\bullet}|^{3/2}$, which allows a straightforward analysis of a wide range of substructure potentials.
The velocity dispersion of stars trapped in a potential $\Phi_{\bullet}$ in steady-state can be found from (4) by solving the isotropic Jeans equations
$\sigma_{\star}^{2}(r)$ $=\frac{1}{\delta_{\star}(r)}\int_{r}^{\infty}{\rm d}r^{\prime}\delta_{\star}(r^{\prime})\bigg{|}\frac{{\rm d}\Phi_{\bullet}}{{\rm d}r}\bigg{|}_{r^{\prime}}$ (5)
$=\frac{2}{5}|\Phi_{\bullet}|$
$=\frac{2}{5}\frac{GM_{\bullet}}{r+c_{\bullet}}.$
which shows that $\sigma_{\star}/v_{e}=(1/5)^{1/2}\approx 0.45$ at all radii. It is important to emphasize that Equation (5) is independent of the speed of the dark object across the host galaxy ($V_{\bullet}$).
Of particular relevance is the distance at which the density of stars equals that of the galactic background. It is convenient to examine the point-mass limit first. Solving $\delta(r_{\epsilon})=1$ in Equation (4) with $c_{\bullet}\to 0$ and $\alpha=1$ yields
@@
For objects with an extended mass distribution, $\delta_{\star}(r_{\epsilon}^{\rm ext})=1$ happens at $r_{\epsilon}^{\rm ext}=r_{\epsilon}-c_{\bullet}=\kappa\,r_{\epsilon}$, where $\kappa\equiv 1-c_{\bullet}/r_{\epsilon}$ is the compactness of the substructure. It is straightforward to show that the condition $\delta_{\star}>1$ requires $\kappa>0$ (or $c_{\bullet}<r_{\epsilon}$), which means that only substructures that are sufficiently compact can contain an over-dense region of captured field stars.
We inspect this issue in next Section with the aid of $N$-body tools.
@@
2.1.2 Strong perturbations
The statistical theory presented above assumes that the substructure potential $\Phi_{\bullet}$ induces weak perturbations on the trajectories of field stars as they move across the volume $V=4\pi r^{3}/3$. One can easily show that this approximation breaks down on scales $r\lesssim r_{\epsilon}$, where the potential well $|\Phi_{\bullet}|$ becomes deeper than the specific kinetic energy of field particles, $K=3\sigma^{2}/2$.
To derive the distribution of captured stars at small distances from the substructure we use two empirical results found in Paper I (see also §3): (i) the phase-space density of captured stars becomes approximately constant in the vicinity of the substructure, such that $f_{\star}({\bf r},{\bf v})\simeq f_{0}$ on scales $r\lesssim r_{\epsilon}$, and (ii) the population of captured stars reaches a steady state on time scales comparable to Smoluchowski’s (1916) ‘fluctuation mean life’ defined by Equation (3).
The density profile of an equilibrium ensemble of stars moving in the potential $\Phi_{\bullet}$ can be derived from the local distribution function without a priori knowledge of their orbital trajectories. That means, there is no need to specify whether orbits are regular (permanently bound) or chaotic (temporarily bound), as long as the distribution of these objects is homogeneous in phase-space. Hence,
$=4\pi f_{0}\,\int_{0}^{|\Phi_{\bullet}|}{\rm d}E[2(E-\Phi_{\bullet})]^{1/2}$
$=\frac{8\sqrt{2}\pi}{3}f_{0}|\Phi_{\bullet}|^{3/2},$
whereas the isotropic 1D velocity dispersion is
$=\frac{2}{5}|\Phi_{\bullet}|.$
Notice that the velocity dispersion profile (9) matches Equation (5), whereas the density profile (8) recovers Equation (4) for a local Maxwellian distribution function
Hence, the profiles derived at distances $r\gtrsim r_{\epsilon}$ under the condition that $\Phi_{\bullet}$ induces weak perturbations also extend to regions where $\Phi_{\bullet}$ dominates the local gravitational field, $r\lesssim r_{\epsilon}$. We have numerically checked that captured particles have a constant phase-space density in dark substructures with an extended mass profile (see Section 3), which agrees with the numerical results shown in Paper I in which the intermediate object is a point-mass.
By construction, combination of Equations (8), (9) and (10) returns a mean phase-space density of bound stars that remains constant across the volume element
Equation (11) highlights two interesting results. First, $Q_{\star}$ solely depends on the potential $\Phi_{\bullet}$ through the value of the abundance parameter, $\alpha$, and second the phase-space density of captured stars drops exponentially for a potential $\Phi_{\bullet}$ that is not at rest with the background ($V_{\bullet}>0$). Numerical tests in §3.1 (see also Appendix 3) discuss the accuracy of these theoretical profiles.
@@ 2.2 Application: Milky Way dwarf spheroidals
In the MW halo, capture of field stars is inefficient on account of their high velocity dispersion. For illustration, here we adopt $\sigma^{\rm MW}\sim 124\,{\rm km\,s^{-1}}$ and $n^{\rm MW}=10^{-4}\,{\rm pc}^{-3}$ at $R=20\,{\rm kpc}$ from the MW centre (e.g. Deason et al. 2011), noting in passing that capture of field stars in this particular stellar halo model becomes systematically less efficient at larger distances. Note that the velocity dispersion of the MW stellar halo is approximately one order of magnitude higher than in the classical dSphs. As a result, the minimum substructure mass needed to capture stars from the field increases up to $M_{\rm min}\gtrsim 10^{8}\,{\rm M}_{\odot}$, which is $\gtrsim 3$ orders of magnitude larger than in the MW dwarf spheroidals. Crucially, this mass threshold is lower than the virial mass of the classical dSphs, which appear to be embedded in DM haloes with virial masses $10^{9}$$10^{10}\,{\rm M}_{\odot}$ (e.g. Peñarrubia et al. 2008; Errani et al. 2018), suggesting that dSphs may contain a population of captured MW halo stars. However, field haloes in this mass range have scale radii that are a factor $\gtrsim 100$ larger than the thermal critical radius plotted in the bottom panel of Fig. 1, $c_{\bullet}\gtrsim 10^{3}\,{\rm pc}$ (e.g. Joyce & Diemand 2019), which implies a negative compactness parameter, $\kappa<0$. Therefore, field stars trapped in dSphs will appear as a diffuse envelope of co-moving stars with subdominant densities with respect to the local MW background ($\delta_{\star}\ll 1$). Therefore, the detection of this stellar population will be challenging, and may require simultaneous modelling of the kinematic and chemical composition of a large sample of stars at the locations of dSphs. This problem will be explored in a separate contribution.
@@ 2.3 Sub-subhalo populations in a CDM framework
It is interesting to extrapolate the virial mass-scale radius relation of CDM field haloes (white dotted line) and subhaloes (white solid line) found in the Aquarius simulations (Springel et al. 2008) down to the relevant mass scales [^6] More precisely, we adopt the power-law fits of the $r_{\rm max}-V_{\rm max}$ measurements plotted in fig. 26 of Springel et al. (2008). As proxy for total halo mass, we use the halo virial mass $M_{200}$ derived from $r_{\rm max}-V_{\rm max}$ under the assumption that field haloes as well as subhaloes follow NFW profiles. For Hernquist spheres the circular velocity peaks at the scale radius, i.e. $r_{\rm max}=c_{\bullet}$.. The first noteworthy result is that subhaloes are more compact than field haloes. This is because substructures that lose a significant amount of mass to tides tend to be more concentrated than field halos of the same mass (e.g. Errani & Navarro 2021). In particular, extrapolation of the Aquarius subhaloes down to $M_{\bullet}\gtrsim 10^{5}\,{\rm M}_{\odot}$ leads to scale-radii $c_{\bullet}\gtrsim 25\,{\rm pc}$, whereas field haloes of the same mass are a factor $\sim 2$ larger, $c_{\bullet}\gtrsim 50\,{\rm pc}$ (Ludlow et al. 2016; see also Diemer & Joyce 2019).

### v2

@@ Abstract
We use analytical and $N$-body methods to study the capture of field stars by gravitating substructures moving across a galactic environment. The majority of stars captured by a substructure move on temporarily-bound orbits that are lost to galactic tides after a few orbital revolutions. In numerical experiments where a substructure model is immersed into a sea of field particles on a circular orbit, we find a population of particles that remain bound to the substructure potential for indefinitely-long times. This population is absent from substructure models initially placed outside the galaxy on an eccentric orbit. We show that gravitational capture is most efficient in dwarf spheroidal galaxies (dSphs) on account of their low velocity dispersions and high stellar phase-space densities. In these galaxies ‘dark’ sub-subhaloes which do not experience in-situ star formation may capture field stars and become visible as stellar overdensities with unusual properties: (i) they would have a large size for their luminosity, (ii) contain stellar populations indistinguishable from the host galaxy, and (iii) exhibit dark matter (DM)-dominated mass-to-light ratios. We discuss the nature of several ‘anomalous’ stellar systems reported as star clusters in the Fornax and Eridanus II dSphs which exhibit some of these characteristics. DM sub-subhaloes with a mass function ${\rm d}N/{\rm d}M_{\bullet}\sim M_{\bullet}^{-\alpha}$ are expected to generate stellar systems with a luminosity function, ${\rm d}N/{\rm d}M_{\star}\sim M_{\star}^{-\beta}$, where $\beta=(2\alpha+1)/3=1.6$ for $\alpha=1.9$. Detecting and characterizing these objects in dSphs would provide unprecedented constraints on the particle mass and cross section of a large range of DM particle candidates.
@@ 2 Statistical theory
Section 2.1 follows the steps outlined in Paper I to compute the average number of field particles that have negative binding energies $E=v^{2}/2+\Phi_{\bullet}(r)<0$ within a spherical volume $V=4\pi r^{3}/3$, where $r$ and $v$ are measured relative to the substructure, as well as their number density ($n_{\star}$) and velocity dispersion ($\sigma_{\star}$) profiles in dynamical equilibrium within a generic potential $\Phi_{\bullet}$. Here, we will focus on capture of field stars, although this analysis can be extended to other tracers moving in the host galaxy potential, including gas particles [^4] To model the dynamics of gaseous particles, a pressure term must be included in the equations of motion presented in §3.1.
@@ 2.1.1 Weak perturbations
The number of field stars temporarily bound to a moving substructure sourcing a potential $\Phi_{\bullet}$ can be estimated statistically from the (local) distribution function under the following simplifying assumptions: (i) field stars move on uncorrelated (random) trajectories within a small volume element $V$, (ii) their number density is roughly constant within the volume $V$, such that $n({\textbf{{R}}}_{\bullet}+{\textbf{{r}}})\approx n({\textbf{{R}}}_{\bullet})=n=N/V$, where ${\textbf{{R}}}_{\bullet}$ is the 3D position vector of the substructure within the host galaxy, and $r=|\mbox{\boldmath$R$}-\mbox{\boldmath$R$}_{\bullet}|$ is the relative distance of a star from the centre of the potential $\Phi_{\bullet}$. This is known as the local approximation, and holds insofar as the (local) density profile rolls slowly, i.e. $r\ll d=|\nabla n/n|^{-1}$. (iii) The relative velocity distribution of particles within the volume element $V=4\pi r^{3}/3$ follows a Maxwellian distribution displaced by the reflex velocity of the point-mass $p(\mbox{\boldmath$v$})=(2\pi\sigma^{2})^{-3/2}\exp[-(\mbox{\boldmath$v$}+\mbox{\boldmath$V$}_{\bullet})^{2}/(2\sigma^{2})$ where $\sigma=\sigma(\mbox{\boldmath$R$}_{\bullet})$ is the local, one-dimensional velocity dispersion of field stars. Under the Maxwellian approximation, the mean-squared (relative) velocity between the background particles and the substructure is $\langle v^{2}\rangle=3\sigma^{2}+V_{\bullet}^{2}$, where ${\textbf{{V}}}_{\bullet}$ is the velocity vector of the substructure with respect to the host galaxy centre.
@@
The approximation that field stars move on uncorrelated trajectories is accurate insofar as the number of stars perturbed by $\Phi_{\bullet}$ represents a small fraction of the total number of stars within the volume element, $N=n\,V$. From Equation (2), it is straightforward to show that the fraction of bound stars within the volume $V$ scales as $N_{b}/N\sim(v_{e}/\sigma)^{3}\,e^{-V_{\bullet}^{2}/(2\sigma^{2})}$ at $r\gtrsim c_{\bullet}$, hence field particles can be treated independently from each other when the escape velocity is low, $v_{e}\lesssim\sigma$, or the substructure velocity is high, $V_{\bullet}\gg\sigma$.
The orbits of field stars captured by the potential $\Phi_{\bullet}$ exhibit chaotic fluctuations of energy and angular momentum, as shown in §3.2. Statistically, one can count how many stars are bound to the substructure potential (i.e. $E<0$) at any given time and compare that number against the value estimated in Equation (2). Hénon & Petit (1986) found that captured particles moving on chaotic orbits only remain trapped within the potential $\Phi_{\bullet}$ over a finite amount of time before being lost to galactic tides. As shown in Paper I, this leads to a population of bound field stars that reaches a steady state as the capture rate (defined as the net number of field particles with $E$ flipping from positive to negative values) equals the loss rate (i.e the net number of field particles with $E$ flipping from negative to positive values). Paper I finds that steady state is typically reached over Smoluchowski’s (1916) ‘fluctuation mean life’
@@
2.1.2 Strong perturbations
The statistical theory presented above assumes that the substructure potential $\Phi_{\bullet}$ induces weak perturbations on the trajectories of field stars as they move across the volume $V=4\pi r^{3}/3$. One can easily show that this approximation breaks down close to the substructure, where the potential well $|\Phi_{\bullet}|$ may become deeper than the specific kinetic energy of field particles, $K=3\sigma^{2}/2$.
To derive the distribution of captured stars at small distances from the substructure we use two empirical results found in Paper I (see also §3): (i) the phase-space density of captured stars becomes approximately constant in the vicinity of the substructure, such that $f_{\star}({\bf r},{\bf v})\simeq f_{0}$, and (ii) the population of captured stars reaches a steady state on time scales comparable to Smoluchowski’s (1916) ‘fluctuation mean life’ defined by Equation (3).
The density profile of an equilibrium ensemble of stars moving in the potential $\Phi_{\bullet}$ can be derived from the local distribution function without a priori knowledge of their orbital trajectories. That means, we do not need to specify whether orbits are regular (permanently bound) or chaotic (temporarily bound), as long as the distribution of these objects is homogeneous in phase-space. Under this condition,
$=4\pi f_{0}\,\int_{0}^{|\Phi_{\bullet}|}{\rm d}E[2(E-\Phi_{\bullet})]^{1/2}$
$=\frac{8\sqrt{2}\pi}{3}f_{0}|\Phi_{\bullet}|^{3/2}.$
To find the normalization of the local phase-space density, $f_{0}$, we match the number density (4) against the steady-state profile derived from Equation (2) at low escape velocities ($v_{e}\lesssim\sigma)$, such that $n_{\star}={\rm d}\,N_{\rm ss}/{\rm d}^{3}r=\alpha\,{\rm d}\,N_{b}/{\rm d}^{3}r$. This yields
which corresponds to a local Maxwellian distribution function multiplied by abundance parameter $\alpha$. The density enhancement induced by the population of energetically-bound stars within the volume $V$ can be found by inserting (5) into (4), which yields
$\delta_{\star}(r)$ $\equiv\frac{n_{\star}(r)}{n}$ (6)
@@ 2.1.2 Strong perturbations
$=\alpha\frac{4}{3\sqrt{\pi}}\frac{(GM_{\bullet})^{3/2}}{\sigma^{3}}e^{-V_{\bullet}^{2}/(2\sigma^{2})}\frac{1}{(r+c_{\bullet})^{3/2}}.$
Notice that the density profile converges to a centrally-divergent ‘density spike’ $\delta_{\star}\sim r^{-3/2}$ (Gondolo & Silk 1999) in the point-mass limit $c_{\bullet}\to 0$. For extended halos ($c_{\bullet}>0$) the profile becomes flat for $r\lesssim c_{\bullet}$. For generic potentials, we find that the density enhancement (6) scales as $\delta_{\star}\sim|\Phi_{\bullet}|^{3/2}$, which allows a straightforward analysis of a wide range of substructure models.
The isotropic 1D velocity dispersion can be derived from the local distribution function as
$=\frac{2}{5}|\Phi_{\bullet}|,$
which is a constant fraction of the escape velocity at all radii, $\sigma_{\star}/v_{e}=(1/5)^{1/2}\approx 0.45$. It is important to emphasize that Equation (7) is independent of the speed of the dark object across the host galaxy ($V_{\bullet}$) as well as of the normalization of the distribution function ($f_{0}$).
By construction, combining Equations (6) and (7) returns a mean phase-space density of bound stars that remains constant across the volume element
Equation (8) highlights two interesting results. First, $Q_{\star}$ solely depends on the potential $\Phi_{\bullet}$ through the value of the abundance parameter, $\alpha$, and second the phase-space density of captured stars drops exponentially for a potential $\Phi_{\bullet}$ that is not at rest with the background ($V_{\bullet}>0$).In §3.1, we carry a number of numerical experiments that inspect the accuracy of the theoretical profiles derived above.
Of particular relevance for this paper is the distance at which the density of bounds stars is equal to that of the field. For reasons that become clear below, it is convenient to examine the point-mass limit first. Solving $\delta(r_{\epsilon})=1$ in Equation (6) with $c_{\bullet}\to 0$ and $\alpha=1$ yields
@@
For objects with an extended mass distribution, $\delta_{\star}(r_{\epsilon}^{\rm ext})=1$ happens at $r_{\epsilon}^{\rm ext}=r_{\epsilon}-c_{\bullet}=\kappa\,r_{\epsilon}$, where $\kappa\equiv 1-c_{\bullet}/r_{\epsilon}$ is the compactness of the substructure. It is straightforward to show that the condition $\delta_{\star}>1$ requires $\kappa>0$ (or $c_{\bullet}<r_{\epsilon}$), which means that only substructures that are sufficiently compact can contain an over-dense region of captured field stars. We inspect this issue in next Section with the aid of $N$-body tools.
@@ 2.2 Application: Milky Way dwarf spheroidals
In the MW halo, capture of field stars is inefficient on account of their high velocity dispersion. For illustration, here we adopt $\sigma^{\rm MW}\sim 124\,{\rm km\,s^{-1}}$ and $n^{\rm MW}=10^{-4}\,{\rm pc}^{-3}$ at $R=20\,{\rm kpc}$ from the MW centre (e.g. Deason et al. 2011), noting in passing that capture of field stars in this particular stellar halo model becomes systematically less efficient at larger distances. Note that the velocity dispersion of the MW stellar halo is approximately one order of magnitude higher than in the classical dSphs. As a result, the minimum substructure mass needed to capture stars from the field increases up to $M_{\rm min}\gtrsim 10^{8}\,{\rm M}_{\odot}$, which is $\gtrsim 3$ orders of magnitude larger than in the MW dwarf spheroidals. Crucially, this mass threshold is lower than the virial mass of the classical dSphs, which appear to be embedded in DM haloes with virial masses $10^{9}$$10^{10}\,{\rm M}_{\odot}$ (e.g. Peñarrubia et al. 2008; Errani et al. 2018), suggesting that dSphs may contain a population of captured MW halo stars. However, field haloes in this mass range have scale radii that are a factor $\gtrsim 100$ larger than the thermal critical radius plotted in the bottom panel of Fig. 1, $c_{\bullet}\gtrsim 10^{3}\,{\rm pc}$ (e.g. Joyce & Diemand 2019), which implies a negative compactness parameter, $\kappa<0$. Therefore, field stars trapped in dSphs will appear as a diffuse envelope of co-moving stars with subdominant densities with respect to the local MW background ($\delta_{\star}\ll 1$). The detection of this stellar population will be challenging, and may require simultaneous modelling of the kinematic and chemical composition of a large sample of stars at the locations of dSphs. This problem will be explored in a separate contribution.